From a group of 6 people, you randomly select 4 of them.

What is the probability that they are the 4 oldest people in the group?

Give your answer as a fraction
Answer #1

Ans:

Total possible ways to choose 4 people from 6 is=6C4

There is only one way to choose the 4 oldest.

P(choosing 4 oldest)=1/6C4=(4/6)*(3/5)*(2/4)*(1/3)

=24/360

=6/90

=1/15